# TIB-STC: A Large-Scale Structured Tibetan Benchmark for Low-Resource Language Modeling

### v4

@@ Abstract
Advancement of large language models (LLMs) has brought transformative capabilities to NLP, but such progress remains unevenly distributed, especially for low-resource and culturally rich languages like Tibetan. In this paper, we present TIB-STC, the first large-scale, expert-curated, and multi-domain benchmark specifically designed to support the development and evaluation of LLMs for the Tibetan language. Spanning over 11 billion tokens across literature, religion, medicine, law, and daily communication, TIB-STC preserves traditional grammar and stylistic richness. To validate its utility, we train a reference model, Sun-Shine, on TIB-STC through a three-stage pipeline involving pretraining, supervised fine-tuning, and preference optimization. Evaluation on TLUE Benchmark for Tibetan-specific tasks, including Ti-MMLU and Ti-SafetyBench, demonstrates the benchmark’s effectiveness in enabling robust instruction-following and culturally aligned generation. We release TIB-STC to advance research in low-resource language modeling and promote inclusivity in multilingual NLP. All data are available: https://github.com/Vicentvankor/sun-shine.
@@ Introduction
Large language models (LLMs) have demonstrated impressive performance across a broad range of natural language processing (NLP) tasks, exhibiting strong generalization and emergent capabilities (Minaee et al. 2024). However, these advancements have primarily benefited high-resource languages such as English and Chinese (Achiam et al. 2023; Yang and et al. 2024; Touvron et al. 2023), while low-resource and linguistically complex languages—such as Tibetan—remain significantly underserved (Liang et al. 2024).
@@
Moreover, mainstream LLMs are typically trained on corpora heavily biased toward dominant languages and cultures (Üstün et al. 2024), limiting their generalization to minority languages. As a result, they often produce outputs that are inaccurate, unsafe, or culturally inappropriate when applied to underrepresented languages (Chao et al. 2024). This underscores the urgent need for high-quality, task-aligned benchmarks that support robust and culturally informed modeling for low-resource languages (Brown et al. 2020; Blodgett et al. 2020).
To address this gap, we introduce TIB-STC, the first large-scale, structured, and expert-curated Tibetan corpus benchmark. Covering over 11 billion tokens from domains such as literature, law, religion, medicine, and daily communication, TIB-STC is designed to support all key stages of LLM development, including pretraining, instruction tuning, safety alignment, and preference optimization. To validate the benchmark’s utility, we also release Sun-Shine, a reference Tibetan LLM based on the LLaMA architecture, trained using a three-stage pipeline. While the primary contribution of this work lies in the construction of the TIB-STC dataset, Sun-Shine serves as a proof of concept for building culturally aligned and linguistically robust LLMs for Tibetan.
@@
  • We introduce TIB-STC, the first large-scale and structured Tibetan corpus benchmark, curated by domain experts across diverse categories including religion, law, medicine, and literature, with over 11B tokens for pretraining and instruction tuning.
  • We have constructed four task-specific sub-benchmarks (Alpaca-Ti, Safety-Prompts-Ti, CValues-Ti, and hh-rlhf-Ti) that support instruction tuning, safety alignment, and preference optimization in low-resource settings.
  • We release Sun-Shine, a reference Tibetan LLM trained on TIB-STC, to demonstrate the benchmark’s effectiveness in instruction-following and safety-aware generation.
@@
Tibetan remains one of the least-resourced languages in NLP due to its complex grammatical structure, two-dimensional syllabic composition, and limited digitized resources (Liu et al. 2024). The scarcity of Tibetan corpora has severely hindered the development of both traditional NLP tools and modern deep learning models (Liu et al. 2025a, b). Existing work in Tibetan NLP has primarily focused on morphological segmentation, part-of-speech tagging, and rule-based translation, often relying on small, domain-specific datasets or handcrafted features (Liang et al. 2024). However, these approaches lack scalability and generalizability, particularly in the context of large-scale pretraining. Moreover, the diverse functional domains of Tibetan, spanning religious scripture, legal and medical texts, and everyday communication, necessitate a corpus that captures both linguistic richness and cultural depth (Khysru et al. 2021). To date, no large-scale, structured benchmark exists to systematically support foundation model training or evaluation for Tibetan.
Our work addresses this gap by introducing TIB-STC, the first comprehensive Tibetan benchmark designed to support all key stages of LLM development, including pretraining, instruction tuning, safety alignment, and preference optimization.
@@
TIB-STC, the first comprehensive and detailed Tibetan corpus benchmark for LLM, consists of five categories of data as shown in the benchmark branch of Table 1. All data selection and cleaning are conducted by experts in the Tibetan language. Unlike modern Tibetan, this benchmark preserves traditional grammar and syntax as much as possible, minimizing the influence of contemporary culture while maintaining its distinct linguistic and cultural characteristics. For the TIB-STC, Corpus is used for LLM pre-training, while other benchmarks focus on optimization and enhancement. For these four sub-benchmarks, Alpaca-Ti and Safety-Prompts-Ti are used for instruction fine-tuning, and CValues-Ti and hh-rlhf-Ti can improve LLM’s alignment ability.
Benchmark Instance Tokens Storage Size
@@
Alpaca-Ti is used for instruction fine-tuning. We draw inspiration from the Alpaca (Taori et al. 2023b, a) project at Stanford University and develop the first Tibetan benchmark specifically designed for self-supervised instruction fine-tuning in the world. It covers the following topics as shown in Table 2.
@@
Safety-Prompts-Ti aims to improve the safety of large models on two major categories: Instruction Attacks and Typical Safety Scenarios and the data size is 38,156. We categorize a total of 15,000 Instruction Attacks into six subtypes, including Goal Hijacking, Prompt Leaking, RolePlay Instructions, Unsafe Instruction Topics, Inquiries with Unsafe Opinions, and Reverse Exposure—each contributing 2,500 instances. In addition, we compile 23,000 samples under Typical Safety Scenarios, covering eight safety-critical topics: Insult (3.6K), Unfairness and Discrimination (3.8K), Crimes and Illegal Activities (3.6K), Physical Harm (1.7K), Mental Health (4K), Privacy and Property (1.3K), Ethics and Morality (1.3K), and Others (4.7K). LLM’s representation space, harmful and harmless queries can be largely distinguished, but this is not noticeably enhanced by safety prompts (Chao et al. 2024; Zheng et al. 2024; Sun et al. 2023). workload it entails, we opted to create a benchmark to simulate the process in advance. One data sample is shown in Figure 3.
@@
CValues-Ti is the benchmark to improve the security of Tibetan language models. Considering work (Moore, Deshpande, and Yang 2024; Rozen et al. 2024; Xu et al. 2023), we propose this branch of a small benchmark because LLM can pose risks or have negative social impacts. For CValues-Ti, it has three categories (safe & responsibility = rejection & positive advice; safe = rejection; unsafe = risk response). Then, under the same prompt, various types of responses can be categorized into positive and negative samples, differentiated by their levels of difficulty, so that there are three common combinations (safe + unsafe; safe & reponsibility + unsafe; safe& reponsibility + safe). Unlike the work (Xu et al. 2023), we streamlined and condensed the instructions, and compared to their benchmark (20K), our sub-benchmark, CValues-Ti (2K), achieved equivalent results in subsequent experiments. One data sample is shown in Figure 4.
@@
hh-rlhf-Ti is also the first benchmark that improves the model effect by introducing human intervention. It consists of two parts: ”harmless (23,822)” and ”helpful (23,589)” just like RLHF (Bai et al. 2022; Kirk et al. 2024), but we reset and summarize the previous benchmark based on the characteristics of Tibetan grammar, and introduce translation, which can either be integrated with or conducted before training for other objectives. This approach could be important in the future to prevent the development of intermediate, unaligned AI systems for Tibetan. It has a total of 47,411 instances, and one data sample is shown in Figure 5.
@@ Conclusion
In this paper, we introduce two key contributions to Tibetan NLP: the TIB-STC benchmark and the Sun-Shine large language model. Our primary focus is the development of TIB-STC, the first large-scale, high-quality Tibetan corpus designed to address the severe lack of resources in low-resource language research. Covering diverse domains such as literature, religion, news, and dialogue, TIB-STC offers structured, culturally rich data that enables the effective training and evaluation of Tibetan language models. Leveraging this benchmark, we further present Sun-Shine, a language model trained on TIB-STC, which demonstrates robust performance in Tibetan-specific tasks, particularly in classical Tibetan understanding and culturally sensitive expression. Extensive evaluation shows that models trained on TIB-STC outperform existing LLMs across metrics such as accuracy, response rate, and stylistic alignment. These results validate the effectiveness of TIB-STC as a foundational resource for Tibetan NLP and highlight the critical role of culturally grounded benchmarks in advancing low-resource language technologies.

### v5

# TIB-STC: A Large-Scale Structured Tibetan Dataset for Low-Resource Language Modeling

@@ Abstract
Advancement of large language models (LLMs) has brought transformative capabilities to NLP, but such progress remains unevenly distributed, especially for low-resource and culturally rich languages like Tibetan. In this paper, we present TIB-STC, the first large-scale, expert-curated, and multi-domain dataset specifically designed to support the development and evaluation of LLMs for the Tibetan language. Spanning over 11 billion tokens across literature, religion, medicine, law, and daily communication, TIB-STC preserves traditional grammar and stylistic richness. To validate its utility, we train a reference model, Sun-Shine, on TIB-STC through a three-stage pipeline involving pretraining, supervised fine-tuning, and preference optimization. Evaluation on TLUE Benchmark for Tibetan-specific tasks, including Ti-MMLU and Ti-SafetyBench, demonstrates the TIB-STC’s effectiveness in enabling robust instruction-following and culturally aligned generation. We release TIB-STC to advance research in low-resource language modeling and promote inclusivity in multilingual NLP. All data are available: https://github.com/Vicentvankor/sun-shine.
@@ Introduction
Large language models (LLMs) have demonstrated impressive performance across a broad range of natural language processing (NLP) tasks, exhibiting strong generalization and emergent capabilities (Minaee et al. 2024). However, these advancements have primarily benefited high-resource languages such as English and Chinese (Achiam et al. 2023; Yang and et al. 2024; Touvron et al. 2023), while low-resource and linguistically complex languages, such as Tibetan, remain significantly underserved (Liang et al. 2024).
@@
Moreover, mainstream LLMs are typically trained on corpora heavily biased toward dominant languages and cultures (Üstün et al. 2024), limiting their generalization to minority languages. As a result, they often produce outputs that are inaccurate, unsafe, or culturally inappropriate when applied to underrepresented languages (Chao et al. 2024). This underscores the urgent need for high-quality, task-aligned datasets that support robust and culturally informed modeling for low-resource languages (Brown et al. 2020; Blodgett et al. 2020).
To address this gap, we introduce TIB-STC, the first large-scale, structured, and expert-curated Tibetan corpus dataset. Covering over 11 billion tokens from domains such as literature, law, religion, medicine, and daily communication, TIB-STC is designed to support all key stages of LLM development, including pretraining, instruction tuning, safety alignment, and preference optimization. To validate the dataset’s utility, we also release Sun-Shine, a reference Tibetan LLM based on the LLaMA architecture, trained using a three-stage pipeline. While the primary contribution of this work lies in the construction of the TIB-STC dataset, Sun-Shine serves as a proof of concept for building culturally aligned and linguistically robust LLMs for Tibetan.
@@
  • We introduce TIB-STC, the first large-scale and structured Tibetan corpus dataset, curated by domain experts across diverse categories including religion, law, medicine, and literature, with over 11B tokens for pretraining and instruction tuning.
  • We have constructed four task-specific sub-datasets (Alpaca-Ti, Safety-Prompts-Ti, CValues-Ti, and hh-rlhf-Ti) that support instruction tuning, safety alignment, and preference optimization in low-resource settings.
  • We release Sun-Shine, a reference Tibetan LLM trained on TIB-STC, to demonstrate the dataset’s effectiveness in instruction-following and safety-aware generation.
@@
Tibetan remains one of the least-resourced languages in NLP due to its complex grammatical structure, two-dimensional syllabic composition, and limited digitized resources (Liu et al. 2024). The scarcity of Tibetan corpora has severely hindered the development of both traditional NLP tools and modern deep learning models (Liu et al. 2025a, b). Existing work in Tibetan NLP has primarily focused on morphological segmentation, part-of-speech tagging, and rule-based translation, often relying on small, domain-specific datasets or handcrafted features (Liang et al. 2024). However, these approaches lack scalability and generalizability, particularly in the context of large-scale pretraining. Moreover, the diverse functional domains of Tibetan, spanning religious scripture, legal and medical texts, and everyday communication, necessitate a corpus that captures both linguistic richness and cultural depth (Khysru et al. 2021). To date, no large-scale, structured dataset exists to systematically support foundation model training or evaluation for Tibetan.
Our work addresses this gap by introducing TIB-STC, the first comprehensive Tibetan dataset designed to support all key stages of LLM development, including pretraining, instruction tuning, safety alignment, and preference optimization.
@@
TIB-STC, the first comprehensive and detailed Tibetan corpus dataset for LLM, consists of five categories of data as shown in the dataset branch of Table 1. All data selection and cleaning are conducted by experts in the Tibetan language. Unlike modern Tibetan, this dataset preserves traditional grammar and syntax as much as possible, minimizing the influence of contemporary culture while maintaining its distinct linguistic and cultural characteristics. For the TIB-STC, Corpus is used for LLM pre-training, while other datasets focus on optimization and enhancement. For these four sub-datasets, Alpaca-Ti and Safety-Prompts-Ti are used for instruction fine-tuning, and CValues-Ti and hh-rlhf-Ti can improve LLM’s alignment ability.
Dataset Instance Tokens Storage Size
@@
Alpaca-Ti is used for instruction fine-tuning. We draw inspiration from the Alpaca (Taori et al. 2023b, a) project at Stanford University and develop the first Tibetan dataset specifically designed for self-supervised instruction fine-tuning in the world. It covers the following topics as shown in Table 2.
@@
Safety-Prompts-Ti aims to improve the safety of large models on two major categories: Instruction Attacks and Typical Safety Scenarios and the data size is 38,156. We categorize a total of 15,000 Instruction Attacks into six subtypes, including Goal Hijacking, Prompt Leaking, RolePlay Instructions, Unsafe Instruction Topics, Inquiries with Unsafe Opinions, and Reverse Exposure—each contributing 2,500 instances. In addition, we compile 23,000 samples under Typical Safety Scenarios, covering eight safety-critical topics: Insult (3.6K), Unfairness and Discrimination (3.8K), Crimes and Illegal Activities (3.6K), Physical Harm (1.7K), Mental Health (4K), Privacy and Property (1.3K), Ethics and Morality (1.3K), and Others (4.7K). LLM’s representation space, harmful and harmless queries can be largely distinguished, but this is not noticeably enhanced by safety prompts (Chao et al. 2024; Zheng et al. 2024; Sun et al. 2023). workload it entails, we opted to create a dataset to simulate the process in advance. One data sample is shown in Figure 3.
@@
CValues-Ti is the dataset to improve the security of Tibetan language models. Considering work (Moore, Deshpande, and Yang 2024; Rozen et al. 2024; Xu et al. 2023), we propose this branch of a small dataset because LLM can pose risks or have negative social impacts. For CValues-Ti, it has three categories (safe & responsibility = rejection & positive advice; safe = rejection; unsafe = risk response). Then, under the same prompt, various types of responses can be categorized into positive and negative samples, differentiated by their levels of difficulty, so that there are three common combinations (safe + unsafe; safe & reponsibility + unsafe; safe& reponsibility + safe). Unlike the work (Xu et al. 2023), we streamlined and condensed the instructions, and compared to their dataset (20K), our sub-dataset, CValues-Ti (2K), achieved equivalent results in subsequent experiments. One data sample is shown in Figure 4.
@@
hh-rlhf-Ti is also the first dataset that improves the model effect by introducing human intervention. It consists of two parts: ”harmless (23,822)” and ”helpful (23,589)” just like RLHF (Bai et al. 2022; Kirk et al. 2024), but we reset and summarize the previous dataset based on the characteristics of Tibetan grammar, and introduce translation, which can either be integrated with or conducted before training for other objectives. This approach could be important in the future to prevent the development of intermediate, unaligned AI systems for Tibetan. It has a total of 47,411 instances, and one data sample is shown in Figure 5.
@@ Conclusion
In this paper, we introduce two key contributions to Tibetan NLP: the TIB-STC dataset and the Sun-Shine large language model. Our primary focus is the development of TIB-STC, the first large-scale, high-quality Tibetan corpus designed to address the severe lack of resources in low-resource language research. Covering diverse domains such as literature, religion, news, and dialogue, TIB-STC offers structured, culturally rich data that enables the effective training and evaluation of Tibetan language models. Leveraging this dataset, we further present Sun-Shine, a language model trained on TIB-STC, which demonstrates robust performance in Tibetan-specific tasks, particularly in classical Tibetan understanding and culturally sensitive expression. Extensive evaluation shows that models trained on TIB-STC outperform existing LLMs across metrics such as accuracy, response rate, and stylistic alignment. These results validate the effectiveness of TIB-STC as a foundational resource for Tibetan NLP and highlight the critical role of culturally grounded datasets in advancing low-resource language technologies.